OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above? y = - x + 30 y = 3x + 6 y = - x + 6 y = x + 30

OpenStudy (texaschic101):

3x + 5y = 30 (solve for y) -- subtract 3x from both sides 5y = -3x + 30 -- divide both sides by 5 y = -3/5x + 30/5 -- reduce when you can y = -3/5x + 6

OpenStudy (anonymous):

y=3x+6 find value os 3x fro this and put it in orignal linear eq..the value os 3x will be 3x=y-6 put it in linear eq. so y-6+5y=30 =>6y=30+6 =>6y=36 =>y=6 answer